### v2

# Bayesian Double Descent

Abstract
Double descent is a phenomenon of over-parameterized statistical models. Our goal is to view double descent from a Bayesian perspective. Over-parameterized models such as deep neural networks have an interesting re-descending property in their risk characteristics. This is a recent phenomenon in machine learning and has been the subject of many studies. As the complexity of the model increases, there is a U-shaped region corresponding to the traditional bias-variance trade-off, but then as the number of parameters equals the number of observations and the model becomes one of interpolation, the risk can become infinite and then, in the over-parameterized region, it re-descends—the double descent effect. We show that this has a natural Bayesian interpretation. Moreover, we show that it is not in conflict with the traditional Occam’s razor that Bayesian models possess, in that they tend to prefer simpler models when possible. We develop comprehensive theoretical foundations including Dawid’s model comparison theory, Dickey-Savage results, and connections to generalized ridge regression and shrinkage methods. We illustrate the approach with examples of Bayesian model selection in neural networks and provide detailed treatments of infinite Gaussian means models and non-parametric regression. Finally, we conclude with directions for future research.
It is vain to do with more what could be done with less. William of Ockham (c. 1285–1347)
Keywords: Bayesian, Double Descent, Interpolation, Neural Network, Model Selection, Occam’s Razor
1 Introduction
Double descent—a phenomenon of over-parameterized models—shows that an estimator’s risk can re-descend as the number of parameters extends past the interpolation limit. Empirically, the double descent effect was initially observed for high-dimensional neural network regression models and the good performance of these models on such tasks as large language models, image processing, and generative AI methods Nareklishvili et al. [2023].
The double descent effect extends the classical bias-variance trade-off curve that shrinkage estimators possess. This phenomenon was first observed in the context of linear regression Belkin et al. [2019], where the authors showed that the test error of the estimator can decrease as the number of parameters increases. Bach [2024] extends these results to stochastic regression models.
Interpolators—estimators that achieve zero training error—were then shown to have attractive properties due to the double descent effect Hastie et al. [2022]. Our goal is to show that Bayesian estimators can also possess a double descent phenomenon. Interpolators such as ReLU neural networks Polson et al. [2017] have increased in popularity with many applications such as traffic flow modeling Polson et al. [2017] and high-frequency trading Dixon et al. [2019], among many others.
Occam’s razor—the favoring of simpler models over complex ones—is a natural feature of Bayesian methods that are based on the weight of evidence (a.k.a. the marginal likelihood of the data). To do this, they penalize models with higher complexity via a correction term as in the Bayesian Information Criterion (BIC). This seems inconsistent with the double descent phenomenon. We show that this is not the case, as even though Bayesian methods shift the posterior towards lower-complexity models, highly parameterized Bayesian models can also have good risk properties due to the conditional prior of parameters given the model. We illustrate this with an application to neural network models.
Double descent has been studied from a frequentest point of view in Belkin et al. [2019], Bach [2024]. The phenomenon of double descent is illustrated in Figure 1. The first part of the curve represents the classical U-shaped bias-variance trade-off. The second part demonstrates the double descent phenomenon, where the test error of the estimator can decrease as the model becomes over-parameterized beyond the interpolation threshold. This phenomenon was later observed in the context of deep learning Nakkiran et al. [2021]. The authors showed that the test error of the estimator can decrease as the number of parameters increases.
Figure 1: Stylized double descent curve showing the classical bias-variance trade-off region (left) and the re-descending risk in the over-parameterized regime (right). The interpolation threshold occurs when the number of parameters equals the number of observations.
To illustrate the double descent phenomenon in a concrete setting, we present a detailed example using polynomial regression with Legendre basis functions. This example demonstrates how the test error can exhibit the characteristic U-shaped curve followed by a re-descent as model complexity increases far beyond the interpolation threshold.
Our demonstration uses a one-dimensional regression problem where we attempt to learn a sinusoidal function $f(x)=\sin(5x)$ from a small dataset of only $n=20$ observations sampled from the interval [1,1][-1,1]. We add Gaussian noise with standard deviation $\sigma=0.3$ to simulate realistic measurement error. The choice of a small sample size is crucial for observing double descent, as it creates a regime where the number of model parameters can substantially exceed the number of observations.
We fit polynomial models of varying degrees $d=1,2,\ldots,50$ using Legendre polynomial basis functions
$\hat{f}(x)=\sum_{j=0}^{d}\hat{\beta}_{j}P_{j}(x),$
where $P_{j}(x)$ denotes the $j$th Legendre polynomial. Legendre polynomials provide a numerically stable orthogonal basis that helps avoid the numerical instabilities associated with standard monomial bases in high-degree polynomial fitting. For each degree $d$, we estimate the coefficients using the Moore-Penrose pseudoinverse, which provides the minimum-norm solution when the system is overdetermined (i.e., when $d>n$).
Figure 2 illustrates how model behavior changes dramatically across different polynomial degrees. The four panels show representative cases that capture the key phases of the double descent phenomenon:
Consider the following four representative cases that illustrate the key phases of the double descent phenomenon.
For degree 1 (underparameterized), the linear model is too simple to capture the oscillatory nature of the underlying sine function. As a result, it exhibits high bias and provides a poor fit to both the training and test data.
At degree 5, we reach the classical optimum. Here, the model complexity is well-matched to the underlying function, striking a balance between bias and variance. The model is sufficiently flexible to capture the main features of the sine function without succumbing to severe overfitting.
When the degree increases to 20, the model reaches the interpolation threshold. At this point, the number of parameters matches the number of training observations, allowing the model to perfectly interpolate the training data. However, this perfect fit comes at a cost: the fitted curve displays wild oscillations between data points, resulting in poor generalization to new data.
Finally, at degree 50, the model is highly over-parameterized, with far more parameters than observations. Surprisingly, in this regime, the model achieves better test performance than at the interpolation threshold. This improvement in generalization, despite the model’s capacity to memorize the training data, is a hallmark of the double descent effect.
Figure 2: Double Descent Phenomenon: Polynomial Regression with Different Degrees
Now, let’s plot the MSE curve. We will plot the test error (blue line) and the training error (red line) for different polynomial degrees from 1 to 50.
Figure 3: Bias-Variance Trade-off: Training and Test MSE vs Model Complexity
The key insight from Figure 3 is the characteristic double descent shape in the test error (blue line). The curve exhibits three distinct phases:
  1. 1.
    Classical Regime: For low degrees ($d<5$), increasing model complexity reduces both bias and test error, following the traditional understanding of the bias-variance tradeoff.
  2. 2.
    Interpolation Crisis: Around the interpolation threshold ($d\approx n=20$), test error peaks dramatically as the model begins to perfectly fit the training data while generalizing poorly.
  3. 3.
    Over-parameterized Regime: For very high degrees ($d>30$), test error decreases again, demonstrating that extreme over-parameterization can lead to improved generalization despite the model’s ability to memorize the training data.
This behavior challenges the conventional wisdom that more parameters necessarily lead to worse generalization. The double descent phenomenon arises from the implicit regularization effects of minimum-norm solutions in over-parameterized settings. When $d>n$, the pseudoinverse solution corresponds to the minimum $\ell_{2}$-norm coefficients among all possible interpolating solutions. This implicit bias toward simpler functions can lead to surprisingly good generalization properties.
While this example uses polynomial regression for clarity, the double descent phenomenon has been observed across a wide range of modern machine learning models, including deep neural networks, random forests, and kernel methods. The implications for practice are significant. Given that model selection is time consuming and computationally expensive, this example shows, that instead of spending time to do model selection to find the ”sweet spot” model with 5-degree polynomial, we just over-parametrise and get a good model for free!
This example serves as a concrete illustration of how classical statistical intuitions about model complexity may not apply in contemporary machine learning settings, particularly when dealing with over-parameterized models that have become increasingly common in practice.
[…]
The rest of this paper is outlined as follows. The next subsection provides connections with the existing literature. Section 2 provides a framework for inference on model complexity and parameters, together with defining Bayes risk, which is an essential part of understanding double descent. Section 3 provides a Bayesian definition of double descent. Section 4 studies the double descent phenomenon in a model for Bayesian interpolation. Section 4.1 provides an application to Bayesian interpolation and neural network regression. Section 5 provides comprehensive coverage of generalized ridge regression and shrinkage methods, including infinite Gaussian means models and non-parametric regression. See also MacKay [1992], who provides a discussion of the weight of evidence and its relationship to model selection and hyperparameter regularization. Finally, Section 6 concludes with directions for future research.
1.1 Connection to Existing Literature
The Bayesian paradigm provides a coherent framework to simultaneously infer parameters and model complexity. Double descent (a.k.a. the re-descending nature of Bayes risk as a function of complexity) can then be studied through a Bayesian lens. The optimal predictive rule is adaptive in the sense that it is a weighted average of individual predictors given complexity.
Our approach shows that double descent is intimately related to the a priori specification of the parameter distribution given complexity. While this acts as an automatic Occam’s razor via the marginal likelihood of model complexity, classical measures such as VC-dimension and asymptotic rates of convergence disregard prior (a.k.a. regularization penalties) and as such miss much of the action in the double descent risk.
MacKay [1992] warns that training and test error rates are likely to be ”uncorrelated”. The double descent property may be partly a mirage and a consequence of a priori regularization specifications. Interpolation of training data is simply a property of a highly non-parametric model. As we demonstrate in this paper, empirical risk bounds that approximate true Bayes risk are of limited use for understanding double descent, as are functional class complexity measures such as VC-dimension, because they disregard the crucial role of prior regularization.
One cannot get away from the fact that the Bayes rule is the optimal processing rule Akaike [1974], and the predictive mean is simply the optimal rule. Rubin [1988] provides an interesting perspective on Bayes robustness; see Polson and Sokolov (2025, in preparation) for further discussion. Stone [1974] points to the inconsistencies in selecting hyperparameters based on cross-validation.
2 Bayesian Model Complexity
Let $D$ denote data. Let $\boldsymbol{\theta}_{M}\in\Theta_{M}$ denote a set of parameters under model $M\in\mathcal{M}$. Let $\boldsymbol{\theta}_{M}=(\theta_{1},\ldots,\theta_{p})$ be the $p$-vector of parameters. The Bayesian approach is straightforward: implement the Bayesian paradigm by executing Bayes rule. This requires the laws of probability and not optimization techniques. The notion of model complexity is no different. Let $\mathcal{M}$ denote the space of models. The Bayesian paradigm simply places probabilities over parameters and models given the data, namely $p(\boldsymbol{\theta}_{M},M|D)$, where $D=(y_{1},\ldots,y_{n})$.
This has a number of decompositions. Bayes theorem calculates the joint posterior over parameters and models given data $D$, namely
$P(\boldsymbol{\theta}_{M},M\mid D)=P(\boldsymbol{\theta}_{M}\mid M,D)P(M\mid D).$
Notice how this factors the posterior into two terms: the conditional posterior over parameters given the model and the posterior over models given data.
Evidence given $M$
The key quantity is the weight of evidence (a.k.a. marginal distribution of the data $D$ given the model $M$), defined by
$p(D|M)=\int_{\Theta_{M}}p(D\mid\boldsymbol{\theta}_{M},M)p(\boldsymbol{\theta}_{M}|M)d\boldsymbol{\theta}_{M}.$
Here $p(D\mid\boldsymbol{\theta}_{M},M)$ is the traditional likelihood function. The key conditional distribution, however, is the specification of the prior over parameters $p(\boldsymbol{\theta}_{M}|M)$. As this is used in the marginalization, it can affect the Bayes risk dramatically. Occam’s razor comes from the fact that this marginalization provides a weight of evidence that favors simpler models over more complex ones.
Models should be simple but not any simpler. (Attributed to Einstein)
This leads to a posterior over models, which is calculated as:
$P(M\mid D)$ $=\dfrac{P(D\mid M)P(M)}{P(D)},$
$P(D\mid M)$ $=\int_{\Theta_{M}}P(D\mid\boldsymbol{\theta}_{M},M)p(\boldsymbol{\theta}_{M}|M)d\boldsymbol{\theta}_{M}.$
Notice that this requires a joint prior specification $p(\boldsymbol{\theta}_{M},M)=p(\boldsymbol{\theta}_{M}|M)p(M)$ over parameters and models. The quantity $p(M|D)$ is the marginal posterior for model complexity given the data. There is an equivalent posterior $p(\boldsymbol{\theta}_{M}|D)$ for the parameters. $p(D\mid M)$ is the evidence of the data $D$ given the complexity (a.k.a. conditional likelihood). The full evidence is
$p(D)=\int p(D|M)p(M)dM.$
This has been used to select the amount of hyperparameter regularization; see, for example, MacKay [1992].
We will see that the prior $p(\boldsymbol{\theta}_{M}|M)$ will lead to an Occam’s razor effect, namely that the marginal distribution will favor simpler models. This will be quantified by the Bayesian Information Criterion (BIC). Importantly, this Occam’s razor effect is not in conflict with the Bayesian double descent phenomenon, which emerges from the marginal posterior of models given data and the conditional prior specification $p(\boldsymbol{\theta}_{M}|M)$.
Posterior for Model Complexity
Let $\mathcal{M}=\{f_{\boldsymbol{\theta}}(x):x\in\mathcal{X}\}$ be a model class parameterized by $\boldsymbol{\theta}$. When necessary, we index the parameters in model $M$ as $\boldsymbol{\theta}_{M}$. Let $y=(y_{1},\ldots,y_{n})$ be a vector of signals. The conditional likelihood under model $M$ with parameters $\boldsymbol{\theta}$ is given by $L(y|\boldsymbol{\theta},M)=\prod_{i=1}^{n}f(y_{i}|\boldsymbol{\theta},M)$. We also allow for the possibility that the data are generated from a model $f_{\text{true}}$ that does not belong to the family of models $f_{\boldsymbol{\theta}}$.
Given a prior measure $\Pi(dF)$ over $\mathcal{F}$, the set of distributions, we can calculate the predictive density:
$f_{n}(y_{n+1}|y_{1},\ldots,y_{n})=\int f(y_{n+1})\Pi_{n}(dF),\quad{\rm where}\;\Pi_{n}(dF)=\frac{\prod_{i=1}^{n}f(y_{i})\Pi(dF)}{\int\prod_{i=1}^{n}f(y_{i})\Pi(dF)}.$
Under the family $f_{\boldsymbol{\theta}}$, we can calculate the parameter posterior as:
$p(\boldsymbol{\theta}|y,M)=\frac{\left(\prod_{i=1}^{n}f(y_{i}|\boldsymbol{\theta},M)\right)p(\boldsymbol{\theta}|M)}{p(y|M)},\quad{\rm where}\;p(y|M)=\int\left(\prod_{i=1}^{n}f(y_{i}|\boldsymbol{\theta},M)\right)p(\boldsymbol{\theta}|M)d\boldsymbol{\theta}.$
Here $p(\boldsymbol{\theta}|M)$ is a prior distribution over parameters for model $M$, and $p(y|M)$ is the marginal likelihood of the data implied by model $M$. This marginal likelihood will be central to our definition of Bayesian double descent.
Asymptotic Carrier of Posterior
The asymptotic behavior of the posterior $p(\boldsymbol{\theta}|y,M)$ is of interest. This is quantified by the asymptotic carrier of the posterior, $\mathcal{C}$, which is defined by:
$\mathcal{C}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\Theta_{M}}\int f_{\text{true}}(y)\log\frac{f_{\text{true}}(y)}{f(y|\boldsymbol{\theta},M)}dy=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\Theta_{M}}\mathrm{KL}(f_{\text{true}}||f_{\cdot|\boldsymbol{\theta},M}).$
That is, the posterior over parameters in the model class $\mathcal{M}$ (for a given $M$) converges to the density $f(\cdot|\boldsymbol{\theta}^{*},M)$ where $\boldsymbol{\theta}^{*}$ minimizes the Kullback-Leibler (KL) divergence between the data generating process $f_{\text{true}}$ and the model class. More precisely, for any neighborhood $\mathcal{A}$ of $\mathcal{C}$, we have:
$\lim_{n\rightarrow\infty}P[\boldsymbol{\theta}\in\mathcal{A}|y_{1},\ldots,y_{n},M]=1\quad\text{a.s. }F_{\text{true}}.$
Since the initial paper of Schwartz [1965], there has been much work on the limiting behavior of the posterior when the true model $f_{\text{true}}$ lies in the class $f_{\boldsymbol{\theta}}$ indexed by models. Berk [1966] provides an interesting result on the posterior when the model is “incorrect” (a.k.a. we only calculate posterior probabilities over models in our given class). Bayes is coherent, and in this sense, we can only calculate relative posterior probabilities. We can simply increase the space of models and re-calculate. Lindley shows the rationality of this approach and terms this process “the art of extending the conversation.” See also the open and closed model framework of Bernardo and Smith [2000]. Convergence in model space to the true model can happen very quickly—at an exponential rate.
Art of Extending the Conversation
Bernardo and Smith [2000] discusses the open and closed model framework. This relates to Peirce’s theory of abduction (rather than induction). Suppose another model $M_{J+1}$ is entertained by the researcher.
We can marginalize in the other direction, over models, to obtain the posterior over the parameter of interest. By extending the conversation, we allow ourselves to imagine the prior over parameters as including $M_{J+1}$, namely:
$p(\boldsymbol{\theta})=\sum_{j=1}^{J+1}p(\boldsymbol{\theta}|M_{j})p(M_{j}).$
This allows us to calculate the posterior of the new model:
$p(M_{J+1}|D)=\frac{p(D|M_{J+1})p(M_{J+1})}{p(D)}.$
Or more simply, the relative posterior probability of $M_{J+1}$ to the set $(M_{1},\ldots,M_{J})$.
Then we can incorporate the new $p(M_{J+1}|y)$ into our analysis. The key point is that the other ratios of model probabilities that we calculated are unchanged!
This is simply the principle of coherence, one of the main benefits of the Bayesian paradigm. The evidence $p(D)$ does change, however, as we have included a new model. See also Bartlett’s paradox Bartlett [1957] and the asymptotic behavior of Bayes factors.
Observing unlikely events shouldn’t necessarily change your view of model space. One philosophical question of interest is whether you can learn just from empirical observation alone and the principles of induction. Diaconis and and Mosteller [1989] develop a theory of coincidences. Fisher, in this regard, remarks:
… for the ’one chance in a million’ will undoubtedly occur with no less and no more than its appropriate frequency, however surprised we may be that it should occur to us. R.A. Fisher (1937, p.16)
There is always Occam’s razor—the marginal likelihood (being an integral) will always favor lower-dimensional models over higher-dimensional ones. There is always a bias-variance trade-off in the construction of Bayesian estimators. We now formalize this with a discussion of BIC.
2.1 BIC
The Bayesian Information Criterion (BIC) is a model selection criterion that penalizes the complexity of the model. It is derived from a Bayesian approach. The BIC is defined as:
$\mathrm{BIC}=\log P(D\mid\hat{\boldsymbol{\theta}}_{k},M_{k})-\frac{k}{2}\log n.$
Here $\hat{\boldsymbol{\theta}}_{k}$ is the maximum likelihood estimate of the $k$ parameters in model $M_{k}$, and $n$ is the sample size. As such, there is a penalty $\frac{k}{2}\log n$ for increasing the dimensionality $k$ of the model under consideration.
To understand the BIC, we need to understand the Bayesian approach to model selection. Specifically, the marginal likelihood of the data under model $M_{k}$ (denoted $M$ for simplicity here) is approximated using Laplace’s method:
$P(D\mid M)=\int P(D\mid\boldsymbol{\theta},M)P(\boldsymbol{\theta}\mid M)d\boldsymbol{\theta}\approx P(D\mid\hat{\boldsymbol{\theta}},M)P(\hat{\boldsymbol{\theta}}\mid M)(2\pi)^{k/2}|\det(\mathbf{H}(\hat{\boldsymbol{\theta}}))|^{-\frac{1}{2}}.$
[…]
$\log P(D\mid M)\approx\log P(D\mid\hat{\boldsymbol{\theta}}_{\text{MLE}},M)-\dfrac{k}{2}\log n,$
which is proportional to the BIC. (Note: The exact definition and derivation of BIC can vary slightly, but this captures the essence). The BIC approximation shows how the Bayesian approach naturally penalizes model complexity through the dimensionality term $\frac{k}{2}\log n$.
The Bayesian approach averages over the posterior distribution of models given data. Suppose that we have a finite list of models $M\in\{M_{1},\ldots,M_{J}\}$. Then we can calculate the posterior over models as:
$p(M_{j}|y)=\frac{p(y|M_{j})p(M_{j})}{\sum_{i=1}^{J}p(y|M_{i})p(M_{i})},\quad{\rm where}\;p(y|M_{j})=\int L_{j}(\boldsymbol{\theta}_{j}|y)p(\boldsymbol{\theta}_{j}|M_{j})d\boldsymbol{\theta}_{j}.$
Laplace’s approximation provides a simple Lindley [1961] illustration of how dimensionality is weighted in the Bayesian paradigm. Hence, BIC is related to a log-posterior approximation. Hence, if prior model probabilities $P(M_{j})$ are uniform, then $P(M_{j}\mid D)\propto P(D\mid M_{j})\approx\exp(\mathrm{BIC}_{j})$.
In a more general case, the evidence (a.k.a. marginal likelihood) for hypotheses (a.k.a. models) $M_{i}$ is calculated as follows:
$P(D\mid M_{i})=\int P(D\mid\boldsymbol{\theta},M_{i})P(\boldsymbol{\theta}\mid M_{i})d\boldsymbol{\theta}.$
Laplace approximation, in the one-dimensional case ($k=1$), yields:
$P(D\mid M_{i})\approx P(D\mid\hat{\boldsymbol{\theta}},M_{i})P(\hat{\boldsymbol{\theta}}\mid M_{i})\sqrt{2\pi}\sigma_{\text{post}}.$
[…]
Generally, in the $k$-dimensional case, we have:
$P(D\mid M_{i})\approx P(D\mid\hat{\boldsymbol{\theta}},M_{i})P(\hat{\boldsymbol{\theta}}\mid M_{i})(2\pi)^{k/2}|\det(-\mathbf{H}(\hat{\boldsymbol{\theta}}))|^{-\frac{1}{2}}.$
Here $\mathbf{H}(\hat{\boldsymbol{\theta}})=\nabla^{2}\log(P(D\mid\hat{\boldsymbol{\theta}},M_{i})P(\hat{\boldsymbol{\theta}}\mid M_{i}))$ is the Hessian of the log-posterior function evaluated at the mode $\hat{\boldsymbol{\theta}}$. As the amount of data collected increases, this Gaussian approximation is expected to become increasingly accurate.
Mackay MacKay [1992] proposes the NIC criterion for selection of neural networks.
The key to double descent is that for $p>n$ the Laplace approximation (a.k.a. BIC) is a poor approximation and in fact the prior $p(\boldsymbol{\theta}|m)$ can affect the Bayes risk. Moreover, the Bayesian folklore is that we should always fit the largest possible model as the following shows.
Bayesian Folklore
Fit a model as big as an elephant (Savage). See MacKay reply to Gelman. von Neumann with $4$ parameters you can fit an elephant. This is the opposite. First as large as possible. Bayesian probability accounting will take care of the rest. This is summarized in the following results.
Theorem:
You can calculate probabilities under the nested simpler model $m$ from the full model $M$. Simply a question of computational power.
This is the Dickey-Savage density ratio: write the parameter vector as $\boldsymbol{\theta}_{M}=(\boldsymbol{\theta}_{m},\boldsymbol{\theta}_{M-m})$ and assume the priors are ”consistent” across specifications. Then we have the result:
$\frac{p(y|m)}{p(y|M)}=\frac{p(\boldsymbol{\theta}_{M-m}=0|y,M)}{p(\boldsymbol{\theta}_{M-m}=0|M)}$
You just need the posterior ordinates for $\boldsymbol{\theta}_{M-m}=0$. Also probabilities calculated under the big model $M$.
The above Bayesian accounting naturally leads to the Dickey-Savage density ratio, which provides a powerful tool for comparing nested models within this framework.
We are assuming that the likelihood for model $m$ is nested in $M$ in the sense that $p(y|\boldsymbol{\theta}_{m})=p(y|\boldsymbol{\theta}_{m},\boldsymbol{\theta}_{M-m}=0)$ and equivalently for the priors:
$p(\boldsymbol{\theta}_{M-m}=0|y)=\int p(\boldsymbol{\theta}_{M-m}=0,\boldsymbol{\theta}_{m}|y)d\boldsymbol{\theta}_{m}=\int\frac{p(y|\boldsymbol{\theta}_{M-m}=0,\boldsymbol{\theta}_{m})p(\boldsymbol{\theta}_{M-m}=0,\boldsymbol{\theta}_{m})}{p(y)}d\boldsymbol{\theta}_{m}$
Assume conditional independence in prior $p(\boldsymbol{\theta}_{M-m}=0,\boldsymbol{\theta}_{m})=p(\boldsymbol{\theta}_{M-m}=0|\boldsymbol{\theta}_{m})p(\boldsymbol{\theta}_{m})$ under $M$, we have
$Var(\boldsymbol{\theta}\mid y)=E\left(Var(\boldsymbol{\theta}\mid m,y)\right)+Var\left(E(\boldsymbol{\theta}\mid m,y)\right).$
Here the pair $\boldsymbol{\theta}_{m},m$ becomes $(f_{m},m)$ and
$f_{m}=\sum_{i=1}^{m}\boldsymbol{\theta}_{j}\phi_{j}(x_{j}).$
Given, $\boldsymbol{\theta}$ is random, we have a random $f$.
Therefore, we have:
$p(\boldsymbol{\theta}_{M-m}=0|D)=\int p(\boldsymbol{\theta}=0,\boldsymbol{\phi}|D)d\boldsymbol{\phi}=\int\frac{p(y|\boldsymbol{\theta}=0,\boldsymbol{\theta}_{m})p(\boldsymbol{\theta}_{m})d\boldsymbol{\theta}_{m}}{p(y)}p(\boldsymbol{\theta}_{M-m}=0)$
Hence, as $p(y|\boldsymbol{\theta}_{M-m}=0)=\int p(y|\boldsymbol{\theta}_{M-m}=0,\boldsymbol{\theta}_{m})p(\boldsymbol{\theta}_{m})d\boldsymbol{\theta}_{m}$, we have:
$\frac{p(\boldsymbol{\theta}_{M-m}=0|y,M)}{p(\boldsymbol{\theta}_{M-m}=0)}=\frac{p(y|\boldsymbol{\theta}_{M-m}=0)}{p(y)}=\frac{p(y|m)}{p(y|M)}$
All calculated under $M$ on the left-hand side!
Hence, we can calculate the posterior over sub-models $m$, purely from $p(\boldsymbol{\theta}_{M}|y,M)$ and hence:
$p(m|y)=\frac{p(\boldsymbol{\theta}_{M-m}=0|y,M)}{p(\boldsymbol{\theta}_{M-m}=0)}p(m)$
2.2 Example: Occam’s Razor
Suppose that we only have two hypotheses $M_{1}$ and $M_{2}$. The posterior odds are given by the prior odds multiplied by the Bayes factor:
$\dfrac{P(M_{1}\mid D)}{P(M_{2}\mid D)}=\underbrace{\dfrac{P(D\mid M_{1})}{P(D\mid M_{2})}}_{\text{Bayes Factor}}\cdot\underbrace{\dfrac{P(M_{1})}{P(M_{2})}}_{\text{Prior Odds}}.$
We can introduce Occam’s razor by putting a higher prior probability $P(M_{1})$ on the simpler model $M_{1}$. The Bayes factor itself automatically embodies Occam’s razor. If $M_{2}$ is a more complex model, it typically spreads its prior probability mass $P(\boldsymbol{\theta}|M_{2})$ over a larger parameter space. This can lead to a smaller marginal likelihood $P(D|M_{2})$ unless the data strongly favor $M_{2}$, even if $M_{2}$ can fit the data perfectly. Thus, if data are fitted about equally well by both models, the simpler model will often have a higher posterior probability, especially if prior odds are equal or favor simplicity.
Example
To illustrate this behavior, suppose that the data are D=(1,3,7,11)D=(-1,3,7,11). Now consider two hypotheses (a.k.a. models) defined by:
  • $M_{a}$: The sequence is an arithmetic progression, $y_{x}=n_{0}+(x-1)n$, where $x=1,2,3,4$, and $n_{0},n$ are integers.
  • $M_{c}$: The sequence is generated by a cubic function of the form $y_{x}=cx^{3}+dx^{2}+ex+f$, where $c,d,e,f$ are fractions. (A cubic can perfectly fit 4 points).
For $M_{a}$, the data $D$ are uniquely generated by $n_{0}=-1,n=4$. Assuming $n_{0},n\in\{-50,-49,\ldots,50\}$ (a space of $101^{2}$ possibilities) and a uniform prior over these, we have:
$P(D\mid M_{a})=\dfrac{1}{101^{2}}\approx 0.000098.$
For $M_{c}$, a cubic polynomial $P(x)=ax^{3}+bx^{2}+cx+d_{0}$ can be made to pass exactly through any four points $(x_{i},y_{i})$. If we assume a very large discrete space of possible fractional coefficients for $c,d,e,f$ such that many combinations exist, but only one (or very few) generates $D$ exactly, then $P(D|M_{c})$ would be (number of fitting parameters) / (total parameter combinations). If the parameter space for $M_{c}$ is much larger than for $M_{a}$, $P(D|M_{c})$ could be smaller than $P(D|M_{a})$, even though $M_{c}$ can also fit the data. (Note: The original example for $M_{c}$ with specific fractions and counts was unclear and potentially flawed; this is a conceptual rewording). Note that we made an implicit assumption that the prior probability on the parameters is uniform within the defined ranges. This assumption is subjective and is not always the appropriate one. For example, if we know that the sequence is generated by a smooth function, we can put a prior on the parameters that is more likely to generate a smooth curve.
Having established how Bayesian methods naturally incorporate Occam’s razor through the marginal likelihood, we now turn to the main contribution of this paper: understanding double descent from a Bayesian perspective.
2.3 Dawid’s Theory of Model Comparison
Let $M\in\mathcal{M}$ and associated with the model $\boldsymbol{\theta}_{M}\in\Theta_{M}$. When model $M$ attains then parameter $\boldsymbol{\theta}_{M}\sim\Pi_{M}(\boldsymbol{\theta}_{M})$ with respect to Lebesgue measure $d\boldsymbol{\theta}_{M}$ over $\Theta_{M}$.
Suppose that data $Y\sim Q$. Then, if $K(Q,M_{2})>K(Q,M_{1})$ where $K$ is the min KL divergence between $Q$ and models in $M$, the log-Bayes factor is given by:
$\log\frac{p(Y|M_{1})}{p(Y|M_{2})}\sim n(KL(Q,M_{2})-KL(Q,M_{1}))+O_{P}(n^{1/2})$
If $Q\in M_{1}\cap M_{2}$, then $KL(Q,M_{2})=KL(Q,M_{1})$ and we have the asymptotic:
$\log\frac{p(Y|M_{1})}{p(Y|M_{2})}=\frac{1}{2}(d_{M_{2}}-d_{M_{1}})\log\left(\frac{n}{2\pi e}\right)+\log\frac{\rho(\boldsymbol{\theta}_{1}^{\star}|M_{2})}{\rho(\boldsymbol{\theta}_{1}^{\star}|M_{1})}+O_{P}(1)$
where $\rho(\boldsymbol{\theta}|M)=\pi_{M}(\boldsymbol{\theta})|\det I_{M}(\boldsymbol{\theta})|^{-1/2}$ is an Occam factor. See also Bartlett (1941) paradox and the asymptotic behavior of Bayes factors.
Having established the Bayesian framework for model complexity and the role of priors in model selection, we now turn to the main contribution of this paper: understanding double descent from a Bayesian perspective. The key insight is that the conditional prior specification $p(\boldsymbol{\theta}_{M}|M)$ acts as an implicit regularizer that can lead to the re-descending risk behavior characteristic of double descent. This connects the theoretical foundations we have developed—including Dawid’s model comparison theory, the Dickey-Savage density ratio, and the BIC approximation—to the empirical phenomenon of double descent.
3 Bayesian Double Descent
3.1 Bayes Risk
The decomposition of Bayes risk in the case of models and parameters leads to an understanding of the phenomenon of double descent. The classic notion of double descent corresponds to the behavior of a risk function $R(M)$ (risk conditional on model complexity $M$), which is defined below. There is a similar decomposition for the overall Bayes risk $R$.
Given an estimator $\hat{\boldsymbol{\theta}}$ of $\boldsymbol{\theta}$, its posterior Bayes risk (expected loss given data $D$) is defined by:
$R(\hat{\boldsymbol{\theta}}|D)=\int L(\hat{\boldsymbol{\theta}},\boldsymbol{\theta})P(\boldsymbol{\theta}\mid D)d\boldsymbol{\theta},$
where $L(\hat{\boldsymbol{\theta}},\boldsymbol{\theta})$ is the loss function. For example, the special case of the squared error loss function is:
$R(\hat{\boldsymbol{\theta}}|D)=\int(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})^{T}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})P(\boldsymbol{\theta}\mid D)d\boldsymbol{\theta}.$
The same framework works in a function space. Given the model $y_{i}=f(x_{i})+\epsilon_{i}$, for a training set $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, the posterior Bayes risk for an estimator $\hat{f}$ of $f$ is:
$R(\hat{f}|D)=\int L(\hat{f},f)P(f\mid D)df.$
For the model complexity problem, we can also consider the prior Bayes risk (or integrated risk), which averages over data $y$, true parameters $\boldsymbol{\theta}$, and models $M$. For squared error loss:
$R_{\text{prior}}(\hat{\boldsymbol{\theta}})=\mathbb{E}_{y,\boldsymbol{\theta},M}\left[(\hat{\boldsymbol{\theta}}(y)-\boldsymbol{\theta})^{2}\right]=\mathbb{E}_{y}\left[\mathbb{E}_{M\mid y}\left[\mathbb{E}_{\boldsymbol{\theta}\mid M,y}\left[(\hat{\boldsymbol{\theta}}(y)-\boldsymbol{\theta})^{2}\right]\right]\right].$
Notice that the optimal estimator (Bayes estimator) under squared error loss is the posterior mean $\hat{\boldsymbol{\theta}}_{\text{Bayes}}(y)=\mathbb{E}\left[\boldsymbol{\theta}\mid y\right]$. By the law of iterated expectations:
$\hat{\boldsymbol{\theta}}_{\text{Bayes}}(y)=\mathbb{E}\left[\boldsymbol{\theta}\mid y\right]=\mathbb{E}_{M\mid y}\left[\mathbb{E}\left[\boldsymbol{\theta}\mid M,y\right]\right].$
Now, we can see a potential “flaw” in interpreting the typical double descent plot (Figure 1) from a purely Bayesian perspective if it only shows $R(\hat{\boldsymbol{\theta}}(M),\boldsymbol{\theta})$ (a frequentist risk for an estimator derived from model $M$). A fully Bayesian approach would average over $p(M\mid y)$. The overall posterior Bayes risk for $\hat{\boldsymbol{\theta}}_{\text{Bayes}}(y)$ can be written as:
$R(\hat{\boldsymbol{\theta}}_{\text{Bayes}}|D)=\mathbb{E}_{M\mid D}\left[R_{M}(\hat{\boldsymbol{\theta}}_{M}|D)\right],$
where $\hat{\boldsymbol{\theta}}_{M}=\mathbb{E}\left[\boldsymbol{\theta}\mid D,M\right]$ and $R_{M}(\hat{\boldsymbol{\theta}}_{M}|D)=\mathbb{E}_{\boldsymbol{\theta}\mid D,M}\left[(\hat{\boldsymbol{\theta}}_{M}-\boldsymbol{\theta})^{2}\right]$.
Definition of Bayesian Double Descent
Let $R(M)=\mathbb{E}_{y,\boldsymbol{\theta}|M}\left[(\hat{\boldsymbol{\theta}}_{M}(y)-\boldsymbol{\theta})^{2}\right]$ be the conditional prior Bayes risk of an estimator $\hat{\boldsymbol{\theta}}_{M}$ derived under model $M$. Let $\hat{\boldsymbol{\theta}}_{M}(y)=\mathbb{E}\left[\boldsymbol{\theta}\mid y,M\right]$ be the optimal Bayes estimator given model $M$. The Bayesian double descent phenomenon occurs when $R(M)$ exhibits a re-descending behavior for $M>n$ (where $M$ represents model complexity and $n$ is the sample size). This re-descent is fundamentally driven by the prior specification $p(\boldsymbol{\theta}_{M}|M)$, which acts as regularization in the over-parameterized regime.
To be more specific, under model $M$
$\hat{f}_{M}=\sum_{i=1}^{M}\hat{\theta}_{i}^{M}\phi_{i}(x)$
Now, we introduce a different smaller model $m$, which is nested in a larger $M$ model
$\hat{f}_{m}=\sum_{i=1}^{m}\hat{\theta}_{i}^{m}\phi_{i}(x)$
and
$\hat{\theta}_{m}=E\left(\boldsymbol{\theta}\mid\boldsymbol{\theta}_{M-m}=,y,M\right).$
Here $\boldsymbol{\theta}_{M-m}$ is the vector
$\boldsymbol{\theta}_{M}=(\boldsymbol{\theta}_{M-m},\boldsymbol{\theta}_{m})$
Training Error and Risk
Consider non-parametric regression, $y_{i}=f(x_{i})+\epsilon_{i}$, where $x_{i}=(x_{1i},\ldots,x_{di})$. We wish to estimate $f(x_{1},\ldots,x_{d})$ where $x=(x_{1},\ldots,x_{d})\in[0,1]^{d}$. From a classical risk perspective, define with $\lVert\cdot\rVert$ denoting the $L^{2}(P_{X})$-norm:
$R(f,\hat{f}_{N})=E_{X,Y}\left[\lVert f-\hat{f}_{N}\rVert^{2}\right].$
Another asymptotic result comes from posterior concentration properties. Let $\hat{f}_{N}$ be constructed as a regularized MAP (maximum a posteriori) estimator which solves the optimization problem:
$\hat{f}_{N}=\operatorname*{arg\,min}_{\tilde{f}_{N}}\left(\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\tilde{f}_{N}(x_{i}))^{2}+\phi(\tilde{f}_{N})\right).$
The ensuing posterior distribution $\Pi(f|x,y)$ can be shown to have optimality results and concentrate at the minimax rate (up to $\log N$ factors). Under standard assumptions, for $\beta$-Hölder smooth functions $f$, we have an optimal minimax rate $\inf_{\hat{f}}\sup_{f}R(f,\hat{f}_{N})$ of $O_{p}(N^{-2\beta/(2\beta+d)})$. Typically, this space is too large as the bound still depends on $d$. By restricting the class of functions, better rates can be obtained, including ones that do not depend on $d$, and in this sense, we avoid the curse of dimensionality. For example, it is common to consider the class of linear superpositions (a.k.a. ridge functions) and projection pursuit.
[…]
$\mathrm{Risk}(\hat{\boldsymbol{\theta}})=\sigma^{2}_{\epsilon}+(\mathbb{E}_{D}[f(x,\hat{\boldsymbol{\theta}}(D))]-f_{0}(x))^{2}+\mathbb{E}_{D}[(f(x,\hat{\boldsymbol{\theta}}(D))-\mathbb{E}_{D}[f(x,\hat{\boldsymbol{\theta}}(D))])^{2}].$
This is (Irreducible Error) + (Squared Bias) + (Variance). (Note: The original text had some confusion in the bias-variance decomposition equations which has been corrected here conceptually).
Given a data set $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, the empirical risk is:
$\hat{R}(\boldsymbol{\theta})=\dfrac{1}{n}\sum_{i=1}^{n}(y_{i}-f(x_{i},\boldsymbol{\theta}))^{2}.$
Minimizing empirical risk is related to maximum likelihood estimation. (Note: The original text’s attempt to decompose empirical risk into bias-variance was problematic and has been omitted here for correctness).
Empirical Risk Bounds
It is worth noting the relationship between expected risk and empirical risk. The expected risk $R(f)$ is typically bounded by the empirical risk $\hat{R}(f)$ plus a complexity term:
$R(f)\leq\hat{R}(f)+O\left(\sqrt{\frac{\text{Complexity}(f)}{N}}\right).$
However, in the case of interpolation, where the model perfectly fits the training data, the empirical risk term $\hat{R}(f)$ becomes zero. This renders such bounds uninformative for interpolators, which are central to understanding double descent. The functional norm or complexity measure is related to concepts like VC-dimension, but the key insight is that classical bounds are often unrelated to prior specifications, which are fundamental to the Bayesian perspective on double descent.
Example: Infinite Gaussian Means
Suppose that we have the infinite Gaussian means problem:
$y_{i}|\boldsymbol{\theta}_{i}$ $\sim N(\boldsymbol{\theta}_{i},\sigma^{2}),$
$\boldsymbol{\theta}_{i}$ $\sim N(0,\tau^{2}).$
This is the well-known variance components ANOVA model. There are many subtle issues in the joint estimation of $(\sigma^{2},\tau^{2})$; see Hill [1965] and Tiao and Tan [1965]. Polson and Scott [2012] provide the global-local shrinkage framework to analyze these models from a fully Bayesian perspective.
There is a long history of studied estimation procedures in these models. For example, Neyman and Scott [1948] shows that the MLE is inconsistent for estimating $\sigma^{2}$. This is due to the fact that one has to marginalize out the infinite-dimensional nuisance parameter $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\ldots)$. Hill [1965] shows that the classical approach can lead to negative variance estimates. This is just a symptom of the fact that there is a spike at the origin in the marginal likelihood for the variance. Hence, the Bayesian paradigm and careful specification of priors are required. Inverse Gamma priors shrink away from the origin and hence do not support the prior assumption of sparsity. Polson and Scott [2012] discuss the use of horseshoe (a.k.a. half-Cauchy scale priors) on the variance components and discuss the analytical behavior of the marginal likelihood.
The heterogeneous variance case is also of interest:
$y_{i}|\boldsymbol{\theta}_{i}$ $\sim N(\boldsymbol{\theta}_{i},\sigma^{2}),$
$\boldsymbol{\theta}_{i}$ $\sim N(0,\tau_{i}^{2}).$
Now we wish to estimate a sequence of prior hyperparameters (a.k.a. regularization parameters) $\mathcal{T}=(\tau_{1}^{2},\tau_{2}^{2},\ldots)$. This will be most useful for the hierarchical model case.
Xu [2007] shows that we need to be careful with the Bayes risk when all the $\tau_{i}$ are zero, as it is infinite. We are expecting non-zero sparsity. A simple estimator for $\tau_{k}^{2}+\sigma^{2}$ is $y_{k}^{2}$ (if $E[\boldsymbol{\theta}_{k}]=0$). We can achieve a better estimator by using empirical Bayes and an adaptive schedule. This requires ordering of the parameters. The Bayes risk can be calculated as follows. The choice of hyperparameter schedule $\{\tau_{i}^{2}\}$ is crucial: if the true $\tau_{k}^{2}$ decay at a certain rate but this rate is misspecified in the prior, it can lead to suboptimal risk behavior and affect the double descent phenomenon.
4 Bayesian Interpolation
We now study the double descent phenomenon in a model for Bayesian interpolation. See Mallick [1998] and Deaton [1980] for a discussion of polynomial regression and related Bayesian approaches. Formally, consider a statistical model of the form:
$y=f(x)+\epsilon,\quad\epsilon\sim N(0,\sigma^{2}).$
Regression then uses a parametric model $f(x,\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ is the parameter vector:
$f(x,\boldsymbol{\theta})=\sum_{j=1}^{q}\theta_{j}\phi_{j}(x),$
where $\phi_{j}$ are basis functions, for example, orthogonal polynomials. A special case of this model is non-parametric regression, when $q$ goes to infinity.
MacKay [1992] discusses Bayesian interpolation. He also discusses the correlation between training and test errors and whether they are related. Specifically, he discusses how the Bayesian paradigm deals with the trade-off between training error and model selection (a.k.a. hyperparameter regularization). The marginal Bayesian probabilities of the model (a.k.a. weight of evidence) provide a more efficient approach and computationally easier methods than traditional classical methods such as cross-validation.
Hierarchical Representation
Young [1977] proposes a Bayesian approach to select polynomial models. He uses a prior that decreases with the degree of $x$. Consider the model:
$y_{i}=h(x_{i})+\epsilon_{i},\quad h(x)=\sum_{j=0}^{q}\theta_{j}p_{j}(x),\quad\mathrm{var}(\epsilon_{i})=\sigma^{2}.$
Here $p_{j}(x)$ are orthogonal polynomials. The likelihood is
$P(D\mid\boldsymbol{\theta},\sigma^{2})=(2\pi\sigma^{2})^{-N/2}\exp\left(-\dfrac{1}{2\sigma^{2}}(y-A\boldsymbol{\theta})^{T}(y-A\boldsymbol{\theta})\right).$
Here $A_{ij}=p_{j}(x_{i})$. The prior is $\boldsymbol{\theta}\sim N(0,C)$, where, for example, $C=\mathrm{diag}(\delta^{2},\tau^{2}/\lambda_{1}^{2},\ldots,\tau^{2}/\lambda_{q}^{2})$. The posterior is $\boldsymbol{\theta}\mid D,\sigma^{2},C\sim N(\hat{\boldsymbol{\theta}}_{\text{post}},\Sigma_{\text{post}})$, where $\Sigma_{\text{post}}=\sigma^{2}(A^{T}A/\sigma^{2}+C^{-1})^{-1}$ and $\hat{\boldsymbol{\theta}}_{\text{post}}=(A^{T}A/\sigma^{2}+C^{-1})^{-1}A^{T}y/\sigma^{2}$.
Efron and Morris [1973] consider the problem of estimating the mean of a multivariate normal distribution. The James-Stein estimator is a form of empirical Bayes estimator. A general Bayes estimator for $\boldsymbol{\theta}$ under normal prior $N(\mu_{0},\tau^{2})$ and normal likelihood $N(\bar{y},\sigma^{2}/n)$ is:
$\hat{\boldsymbol{\theta}}_{\text{Bayes}}=\frac{n\tau^{2}}{\sigma^{2}+n\tau^{2}}\bar{y}+\frac{\sigma^{2}}{\sigma^{2}+n\tau^{2}}\mu_{0}.$
4.1 Neural Network Regression
Deaton [1980] considers the polynomial fit
$Y_{i}=P(x_{i})+\epsilon_{i},\quad\epsilon_{i}\sim N(0,\sigma^{2}).$
Here $P(x)=\sum_{j=0}^{m}\boldsymbol{\theta}_{j}\psi_{j}(x)$. Polynomials $\psi_{j}(x)$ are orthogonal with respect to the data points $x_{k}$, such that $\sum_{k=1}^{N}\psi_{i}(x_{k})\psi_{j}(x_{k})=\delta_{ij}$ (assuming appropriate normalization, so $Q^{T}Q=I$ where $Q_{ki}=\psi_{i}(x_{k})$).
The prior specification is given by a generalized ridge specification: $\boldsymbol{\theta}_{j}\sim N(0,\sigma_{j}^{2})$. The likelihood $Y\mid\boldsymbol{\theta}\sim N(Q\boldsymbol{\theta},\sigma^{2}I)$ is
$P(Y\mid\boldsymbol{\theta},\sigma^{2})=(2\pi\sigma^{2})^{-N/2}\exp\left(-\dfrac{1}{2\sigma^{2}}(Y-Q\boldsymbol{\theta})^{T}(Y-Q\boldsymbol{\theta})\right).$
Given $Q^{T}Q=I$, the OLS estimates are $\hat{\boldsymbol{\theta}}_{\text{OLS}}=Q^{T}Y$, and $\mathrm{var}(\hat{\boldsymbol{\theta}}_{\text{OLS}})=\sigma^{2}I$. Thus, $\hat{\boldsymbol{\theta}}_{i}\mid\boldsymbol{\theta}_{i},\sigma^{2}\sim N(\boldsymbol{\theta}_{i},\sigma^{2})$. If $s=(Y-Q\hat{\boldsymbol{\theta}}_{\text{OLS}})^{T}(Y-Q\hat{\boldsymbol{\theta}}_{\text{OLS}})$ is the residual sum of squares, then $s/\sigma^{2}\sim\chi^{2}_{N-(p+1)}$. Using the Bayes rule for the normal-normal model (prior $\boldsymbol{\theta}_{i}\sim N(0,\sigma_{i}^{2})$, likelihood from $\hat{\boldsymbol{\theta}}_{i}$), the posterior for $\boldsymbol{\theta}_{i}$ is:
$\boldsymbol{\theta}_{i}\mid\hat{\boldsymbol{\theta}}_{i},\sigma^{2},\sigma_{i}^{2}\sim N\left(\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}+\sigma^{2}}\hat{\boldsymbol{\theta}}_{i},\frac{\sigma_{i}^{2}\sigma^{2}}{\sigma_{i}^{2}+\sigma^{2}}\right).$
Let $z_{i}=\dfrac{\sigma^{2}}{\sigma_{i}^{2}+\sigma^{2}}$. Then the posterior mean is $(1-z_{i})\hat{\boldsymbol{\theta}}_{i}$ and posterior variance is $(1-z_{i})\sigma^{2}=z_{i}\sigma_{i}^{2}$. The marginal distribution for $\hat{\boldsymbol{\theta}}_{i}$ (integrating out $\boldsymbol{\theta}_{i}$) is
$\hat{\boldsymbol{\theta}}_{i}\mid\sigma^{2},\sigma_{i}^{2}\sim N(0,\sigma^{2}+\sigma_{i}^{2})=N(0,\sigma^{2}/z_{i}).$
Given d=N(p+1)d=N-(p+1), the joint marginal likelihood of hyperparameters $(\{z_{i}\}_{i=0}^{m},\sigma^{2})$ given data $(\{\hat{\boldsymbol{\theta}}_{i}\}_{i=0}^{p},s)$ is proportional to (from Deaton, 1980, Eq. 2.6, adapted):
$L(\{z_{i}\},\sigma^{2}\mid\hat{\boldsymbol{\theta}},s)\propto(\sigma^{2})^{-N/2}s^{(d-2)/2}\exp\left(-\dfrac{s}{2\sigma^{2}}\right)\prod_{i=0}^{p}z_{i}^{1/2}\exp\left(-\dfrac{\hat{\boldsymbol{\theta}}_{i}^{2}z_{i}}{2\sigma^{2}}\right).$ (1)
Now, the question is how to estimate $z_{i}$ (and $\sigma^{2}$) from data. It is not assumed that $z_{i}$ are all equal. Deaton assumes an ordering: $\sigma_{0}^{2}\geq\sigma_{1}^{2}\geq\ldots\geq\sigma_{p}^{2}$, which implies $0<z_{0}\leq z_{1}\leq\ldots\leq z_{p}\leq 1$. Deaton reparameterizes $V_{i}=z_{i}/\sigma^{2}=1/(\sigma_{i}^{2}+\sigma^{2})$ for $i=0,\dots,p$ and $V_{p+1}=1/\sigma^{2}$. The constraint becomes $V_{0}\leq V_{1}\leq\dots\leq V_{p}\leq V_{p+1}$. The log-likelihood in terms of $V_{i}$ (from Deaton, Eq. 2.8, adapted, assuming $N$ for $n$ in original text’s formula):
$\log L=\text{const}+\frac{d-2}{2}\log s+\frac{N-(p+1)}{2}\log V_{p+1}-\frac{1}{2}V_{p+1}s+\sum_{i=0}^{p}\left(\frac{1}{2}\log\frac{V_{i}}{V_{p+1}}-\frac{1}{2}\hat{\boldsymbol{\theta}}_{i}^{2}V_{i}\right).$
Deaton then adds priors for $V_{i}$. For example, Gamma priors $p(V_{i})\propto V_{i}^{\gamma_{i}-1}e^{-V_{i}/\beta_{i}}$. The unconstrained MAP estimates for $V_{i}$ (from differentiating the log-posterior) are of the form $\hat{V}_{i}^{\text{unc}}=(\text{shape}_{i}-1)/(\text{scale}_{i}+\hat{\boldsymbol{\theta}}_{i}^{2}/2)$. These unconstrained estimates are then adjusted using isotonic regression (e.g., PAVA) to satisfy the order constraint $V_{0}\leq\dots\leq V_{p}\leq V_{p+1}$. The PAVA formula is $V_{j}^{*}=\max_{k\leq j}\min_{l\geq j}\mathrm{Av}(k,l)$, where $\mathrm{Av}(k,l)=\frac{\sum_{r=k}^{l}w_{r}\hat{V}_{r}^{\text{unc}}}{\sum_{r=k}^{l}w_{r}}$ with appropriate weights $w_{r}$. The condition for positive estimates often involves shape parameters $\gamma_{i}>1/2$.
Estimation of $\sigma^{2}$
Deaton [1980] provides an empirical Bayes approach where $V_{p+1}=1/\sigma^{2}$ is estimated as part of the isotonic procedure.
To illustrate Deaton’s approach, consider the case when the true model is as follows. We use normalized Legendre polynomials as the basis:
$y(x,\boldsymbol{\theta})=f(x,\boldsymbol{\theta})+\epsilon,\quad\epsilon_{i}\sim N(0,0.3^{2}).$
In matrix form, for a vector of inputs $\mathbf{x}=[x_{1},x_{2},\ldots,x_{N}]^{T}$, with $N=20$:
$\mathbf{y}=\mathbf{Q}\boldsymbol{\theta}+\boldsymbol{\epsilon}.$
Figure 4: Data generated from the model $y(x,\boldsymbol{\theta})=f(x,\boldsymbol{\theta})+\epsilon$, where $\epsilon_{i}\sim N(0,0.3^{2})$. The red dots represent the true function $f(x,\boldsymbol{\theta})$, and the blue line shows the noisy observations.
Figure 5 shows the marginal likelihood on a logarithmic scale for different model complexities $m$.
Figure 5: Marginal likelihood for a model with true polynomial degree $p_{\text{true}}=10$ (example) and $N=20$ observations. The x-axis represents the assumed model complexity $p$ (degree of polynomial fit), and the y-axis shows the log marginal likelihood. The peak indicates the optimal model complexity as selected by the marginal likelihood.
Having established the Bayesian framework for polynomial regression and neural networks, we now explore how these ideas can be extended to more flexible basis functions that can adapt to the data structure.
4.2 Extensions to Neural Network Basis Functions
The polynomial basis approach illustrated in the previous section can be naturally extended to neural network basis functions. Instead of using fixed orthogonal polynomials like Legendre polynomials, we can employ neural networks to learn adaptive orthogonal basis functions that are tailored to the specific regression problem.
Orthogonal Neural Network Architecture
Consider a neural network architecture that outputs $M$-dimensional vectors, where $M$ is the number of basis functions. The key requirement is that these output vectors must be orthogonal with respect to the input vectors, meaning that the inner product of any two output vectors for different inputs should be zero. This can be achieved through several approaches:
  1. 1.
    Gram-Schmidt Orthogonalization: Apply the Gram-Schmidt process during the forward pass to ensure orthogonality of the basis functions across the training samples.
  2. 2.
    QR Decomposition: Use QR decomposition to orthogonalize the neural network outputs, which provides better numerical stability for larger basis sets.
  3. 3.
    Regularization-Based Orthogonality: Add an orthogonality penalty term to the loss function that encourages the basis functions to be orthogonal:
    $\mathcal{L}_{\text{ortho}}=\lambda\|\mathbf{Q}^{T}\mathbf{Q}-\mathbf{I}\|_{F}^{2}$
    where $\mathbf{Q}$ contains the basis function outputs and $\lambda$ controls the strength of the orthogonality constraint.
  4. 4.
    Learnable Orthogonal Basis: Design separate networks for each basis function and enforce orthogonality through architectural constraints or post-processing.
Bayesian Treatment of Neural Network Basis
The Bayesian framework naturally extends to the neural network case. Let $\phi_{j}(x;\mathbf{w})$ denote the $j$-th neural network basis function with parameters $\mathbf{w}$. The regression model becomes:
$y_{i}=\sum_{j=1}^{M}\boldsymbol{\theta}_{j}\phi_{j}(x_{i};\mathbf{w})+\epsilon_{i},\quad\epsilon_{i}\sim N(0,\sigma^{2})$
The hierarchical Bayesian specification includes priors on both the basis function parameters $\mathbf{w}$ and the linear combination coefficients $\boldsymbol{\theta}$:
$\boldsymbol{\theta}_{j}$ $\sim N(0,\sigma_{j}^{2}),\quad j=1,\ldots,M$ (2)
$\mathbf{w}$ $\sim p(\mathbf{w})$ (3)
$\sigma_{j}^{2}$ $\sim p(\sigma_{j}^{2})$ (4)
This allows for adaptive learning of basis functions while maintaining the Bayesian model selection framework. The marginal likelihood can still be computed (approximately) to select the optimal number of basis functions $M$, potentially exhibiting the double descent phenomenon as $M$ varies.
Computational Considerations
Extending the Bayesian framework to neural network basis functions introduces several computational challenges. Maintaining orthogonality among the learned basis functions during training requires additional mechanisms, such as projection methods or the inclusion of penalty terms in the loss function. Furthermore, the computation of the marginal likelihood becomes substantially more complex, owing to the nonlinear dependence on the neural network parameters $\mathbf{w}$. As a result, standard analytical approaches are often infeasible, and approximate inference techniques—such as variational inference or Markov chain Monte Carlo sampling—may be necessary to perform posterior computations in this setting.
Despite these challenges, neural network basis functions offer several notable advantages over fixed polynomial bases. Most importantly, they provide adaptivity, allowing the basis functions to conform to the specific structure present in the data. Neural networks also possess greater representational power, enabling them to capture more complex patterns than traditional polynomial expansions. In addition, neural networks can automatically learn relevant features from the data, eliminating the need to pre-specify polynomial orders or other basis parameters. Finally, this approach is potentially more scalable, as neural networks are well-suited to handling high-dimensional input spaces where polynomial expansions would become unwieldy.
This neural network extension provides a promising direction for combining the interpretability and theoretical guarantees of Bayesian polynomial regression with the representational power of deep learning, while maintaining the double descent framework we have developed.
The connection between neural networks and ridge regression becomes clearer when we consider the canonical basis representation. In the following section, we develop the theoretical foundations of generalized ridge regression and show how it relates to the Bayesian double descent phenomenon we have been studying.
5 Generalized Ridge Regression and Shrinkage Methods
5.1 Generalized Ridge Regression in the Canonical Basis
In global-local shrinkage models and generalized ridge regression (GRR), the strength of regularization is governed by local penalties $k_{i}$, which correspond to prior precision parameters in the canonical regression basis. Selecting appropriate values for these $\kappa_{i}$’s is essential to achieve optimal bias-variance tradeoffs.
A common approach is to use marginal maximum likelihood or Type II maximum likelihood estimation. Suppose we observe rotated data $z_{i}=d_{i}\hat{\alpha}_{i}$, with marginal distribution:
$p(z_{i}\mid k_{i},\sigma^{2})\sim N\left(0,\sigma^{2}\left(1+\frac{d_{i}^{2}}{\kappa_{i}}\right)\right).$
Let $\phi_{i}^{2}=\sigma^{2}(1+d_{i}^{2}/\kappa_{i})$. By re-parametrizing and maximizing the likelihood with respect to $\kappa_{i}$, we obtain the empirical Bayes estimate:
$\hat{\kappa}_{i}=\frac{d_{i}^{2}\sigma^{2}}{z_{i}^{2}-\sigma^{2}},\quad\text{provided that }z_{i}^{2}>\sigma^{2}.$
This yields a data-adaptive shrinkage rule; coordinates with weak signal ($z_{i}^{2}\leq\sigma^{2}$) receive effectively infinite penalty ($\kappa_{i}=\infty$), corresponding to complete shrinkage. Stronger signals receive smaller penalties and are retained.
Remark 5.1.
This thresholding rule mimics the behavior of the non-negative garotte estimator proposed by breiman1995better, wherein coordinates are retained or eliminated based on their marginal evidence. It also aligns with modern empirical Bayes variable selection ideas.
We revisit the ridge regression framework from the perspective of the canonical (rotated) coordinate system, consistent with the singular value decomposition $X=UDW^{\top}$ and rotated coefficients $\boldsymbol{\alpha}=W^{\top}\boldsymbol{\beta}$. In this basis, the regression model becomes:
$Y=X\boldsymbol{\beta}+\epsilon=UD\boldsymbol{\alpha}+\epsilon,$
and projecting onto $U$, we obtain:
$\hat{\boldsymbol{\alpha}}_{i}=\boldsymbol{\alpha}_{i}+\epsilon_{i},\quad\epsilon_{i}\sim N(0,\sigma^{2}/d_{i}^{2}),$
where $d_{i}$ is the $i$th singular value and $\hat{\boldsymbol{\alpha}}_{i}$ is the least-squares estimate of $\boldsymbol{\alpha}_{i}$.
Now suppose we place independent priors $\boldsymbol{\alpha}_{i}\sim N(0,v_{i})$ on the coefficients in this rotated basis. The posterior mean under this Gaussian model yields a shrinkage estimator of the form:
$\boldsymbol{\alpha}_{i}^{*}=\kappa_{i}\hat{\boldsymbol{\alpha}}_{i},\quad\text{with}\quad\kappa_{i}=\frac{d_{i}^{2}v_{i}}{\sigma^{2}+d_{i}^{2}v_{i}}=\frac{d_{i}^{2}}{d_{i}^{2}+k_{i}},$ (5)
where $k_{i}=\sigma^{2}/v_{i}$ is the generalized ridge penalty for the $i$th component. When all $k_{i}=k$, this recovers standard ridge regression.
Goldstein and Smith [goldstein1974ridgetype] provide conditions under which this class of estimators dominates the least-squares solution in componentwise mean squared error (MSE). Specifically, if the smallest singular value $d_{\min}^{2}$ exceeds a threshold determined by the signal-to-noise ratio, then:
$\text{MSE}(\boldsymbol{\alpha}_{i}^{*})<\text{MSE}(\hat{\boldsymbol{\alpha}}_{i})$
for all $i$. This contrasts with James–Stein estimators, which guarantee overall MSE reduction but not improvement for each parameter individually.
While the James–Stein estimator also shrinks the OLS coefficients toward zero, it does so uniformly across all coordinates through a global multiplier that depends on the total squared norm $\|\hat{\boldsymbol{\alpha}}\|^{2}$. In contrast, the GRR estimator performs shrinkage independently along each coordinate, based on local information—specifically, the corresponding singular value $d_{i}$ and penalty $k_{i}$. This local, componentwise shrinkage anticipates the structure of modern global-local shrinkage priors, where each parameter receives its own adaptive regularization governed by local scale terms. In both settings, the goal is to suppress estimation error more aggressively in directions of low signal-to-noise, but GRR achieves this in a fully decoupled fashion, which enables componentwise risk improvements that the James–Stein estimator cannot guarantee.
Furthermore, one can evaluate predictive performance using the expected predictive mean squared error (PMSE), which for shrinkage weights $\kappa_{i}=\kappa(d_{i},k_{i})$ becomes:
$\text{PMSE}(\tilde{\boldsymbol{\beta}})=\sum_{i=1}^{p}\int\left\{\kappa_{i}z_{i}-\mu_{i}\right\}^{2}dF_{\mu_{i}}(z_{i}),$ (6)
where $z_{i}|\mu_{i}\sim N(\mu_{i},1)$, and $\mu_{i}$ represents the true canonical coefficient.
Thus, generalized ridge regression offers a flexible, interpretable alternative to uniform shrinkage methods, with clear Bayesian and risk-based justification. This framework also accommodates more flexible priors on $k_{i}$, such as inverse-gamma or half-Cauchy, yielding global-local priors with adaptive shrinkage [Carvalho et al., 2010, Polson and Scott, 2012].
5.2 Shrinkage in Infinite Gaussian Means Models
One observes input-output pairs $(y_{i},x_{i})_{i=1}^{N}$ from the non-parametric regression $y_{i}=f(x_{i})+\epsilon_{i}$ where $\epsilon_{i}\sim N(0,\sigma^{2})$. Here $f$ is only assumed to lie in some infinite dimensional smooth class $\mathcal{F}$. The infinite sequence models $Z_{n}=\{Z_{jn},j=1,2,\ldots\}$ where $Z_{jn}\sim N\left(\boldsymbol{\theta}_{j},\frac{\sigma^{2}}{n}\right)$. Let the parameter space be $\Theta=\{\boldsymbol{\theta}_{j}:j=1,2,\ldots\}\subset\mathbb{R}^{\infty}$ a (large) subset.
Here $\boldsymbol{\theta}_{j}=\int_{0}^{1}f(t)\phi_{j}(t)dt$ given an orthonormal basis $\{\phi_{j}\}$ of $L_{2}[0,1]$.
Infinite Gaussian Bayes Risk
Asymptotics of fixed dimension are greatly different for finite (fixed) and infinite cases. To explore the asymptotics in $n$, it is useful to index the parameter as $\boldsymbol{\theta}_{in}$ and to scale accordingly. Let $X=(X_{1n},X_{2n},\ldots)$ and $\hat{\boldsymbol{\theta}}_{n}=(\hat{\boldsymbol{\theta}}_{1n},\hat{\boldsymbol{\theta}}_{2n},\ldots)$ be independent and $\epsilon_{in}\sim N\left(0,\frac{v^{2}}{n}\right)$.
The Bayes estimate $\hat{\boldsymbol{\theta}}_{in}$ of $\boldsymbol{\theta}_{in}$ is:
$\hat{\boldsymbol{\theta}}_{in}=\frac{\tau^{2}_{i}}{\tau_{i}^{2}+\sigma^{2}/n}X_{in}$
The associated Bayes Risk is:
$R_{n}=\frac{\sigma^{2}}{n}\sum_{i=1}^{\infty}\frac{\tau^{2}_{i}}{\tau_{i}^{2}+\sigma^{2}/n}$
Now if we assume a fixed sequence for the prior scales $\boldsymbol{\theta}_{i}|\tau_{i}^{2}\sim N(0,\tau_{i}^{2})$ we can have non-robust asymptotics.
By taking a fixed schedule (power or exponential) of $\tau_{i}$’s you can get different effects. Rubin shows that $R_{n}$ is $O(\log n/n)$ for $\tau_{i}=2^{-i}$ but is $O(1/\sqrt{n})$ for $\tau_{i}=i^{-2}$.
We now consider the problem of estimating a sequence of normal means in a hierarchical setting. This formulation captures the structure of both the James–Stein problem and empirical Bayes estimation under global-local shrinkage.
Suppose the observations follow the hierarchical model:
$y_{i}$ $=\boldsymbol{\theta}_{i}+\epsilon_{i},\quad\epsilon_{i}\sim N(0,\sigma^{2}),$
$\boldsymbol{\theta}_{i}$ $=\boldsymbol{\mu}+\boldsymbol{\tau}u_{i},\quad u_{i}\sim N(0,\mathbf{I}),$
with possibly heterogeneous variances if we allow $\boldsymbol{\tau}=(\tau_{1},\tau_{2},\ldots)$ to vary across components. This corresponds to a classical variance components model, sometimes referred to as the one-way random effects or ANOVA model.
The classical estimation of $(\sigma^{2},\boldsymbol{\tau}^{2})$ is known to be challenging. In particular, the Neyman–Scott paradox [Neyman and Scott, 1948] shows that the MLE of $\sigma^{2}$ is inconsistent when the number of parameters grows with the sample size. In small-sample settings, standard estimators may yield negative variance estimates, as shown by Hill [1965]. These pathologies motivate Bayesian shrinkage methods.
A fully Bayesian approach imposes priors on the variance components. When $\tau_{i}$ is allowed to vary across $i$, this leads to a global-local model. We aim to estimate a sequence $\mathcal{T}=(\tau_{1},\tau_{2},\ldots)$ using the observed data. Rubin (1987) shows that different decay schedules for $\tau_{k}$ lead to dramatically different rates of posterior contraction. For instance:
  • $\tau_{k}=k^{-2}$ yields Bayes risk $\mathcal{O}(n^{-1/2})$
  • $\tau_{k}=2^{-k}$ yields risk $\mathcal{O}(1/\log n)$
  • If the true signal follows $\tau_{k}=2^{-k}$ but prior is $k^{-2}$, then error is $\mathcal{O}(\log n/n)$
These results highlight the importance of adaptively estimating local variance scales, as is done in global-local priors.
Remark 5.2.
The optimal local variance $\tau_{i}^{2}$ can be estimated using the least concave majorant (LCM) of the CUSUM of $X_{in}^{2}-\sigma^{2}/n$, which serves as a form of isotonic regression. This procedure was explored by Rubin and extended by Deaton [Deaton, 1980] for polynomial regression under order constraints.
Xu (2007) showed how least concave majorants (LCM) can be used to estimate a sequence of prior variances $\tau_{i}^{2}$ from observed data $X_{i}^{2}$. The method constructs the LCM of the cumulative sum of squared signals, effectively performing an adaptive isotonic regression on the variance sequence. This procedure recovers structure in sparse or smoothly decaying sequences, and is particularly useful for modeling high-dimensional global-local priors.
Figure 6 reproduces two illustrations from Xu’s thesis. The left panel (Example 2.3) shows the log-CUSUM of $X_{i}^{2}$ (dots) for $r=1,\ldots,1005$, along with its least concave majorant (blue line), providing a smoothed approximation of signal accumulation. The right panel (Example 2.4) compares three quantities on a log-scale: (i) true variance decay $\tau_{i}^{2}=1/i^{2}$ (dashed), (ii) noisy variance proxies $X_{i}^{2}$ (dots), and (iii) refined estimates from the slope of the LCM (blue line). The LCM estimates exhibit superior stability in estimating $\tau_{i}^{2}$.
(a) CUSUM of $X_{i}^{2}$ (dots) and LCM (blue line)
(b) True vs. empirical vs. LCM-refined $\tau_{i}^{2}$ (log scale)
Figure 6: Illustration of Xu’s variance component estimation method via least concave majorants (LCM). Adapted from Figures 2.5 and 2.6 of Xu [2007].
LCM Estimation Formula:
Let $\eta_{r}=\sum_{i=1}^{r}X_{i}^{2}$ and consider the piecewise-linear interpolation of its LCM. Define the slope at each $i$ as:
$\hat{\tau}_{i}^{2}=\left(\min_{1\leq h\leq i}\max_{i\leq j\leq K}\frac{1}{j-h+1}\sum_{l=h}^{j}\left(X_{l}^{2}-\frac{\sigma^{2}}{n}\right)\right)_{+}.$
This isotonic estimator shrinks noisy $X_{i}^{2}$ estimates toward the underlying monotone structure implied by the decay of the true signal.
6 Discussion
This paper has provided a comprehensive Bayesian perspective on the double descent phenomenon. We have shown that double descent is naturally understood from a Bayesian viewpoint, where there is a natural bias-variance trade-off built into the Bayesian paradigm. This also leads to an automatic Occam’s razor—the automatic penalization of model complexity via the marginal likelihood.
The Bayesian paradigm provides a coherent framework to simultaneously infer parameters and model complexity, and thus to understand double descent, namely the re-descending property of Bayes risk as a function of complexity. The optimal predictive rule is adaptive and is a weighted average of individual predictors given complexity.
In this work, we have made several key contributions to the understanding of double descent from a Bayesian perspective. First, we have shown that the conditional prior specification $p(\boldsymbol{\theta}_{M}|M)$ serves as an implicit regularizer, fundamentally driving the double descent phenomenon. This insight provides a natural Bayesian interpretation of double descent, filling a gap that previously existed in the literature.
Second, we have clarified the relationship between Bayesian Occam’s razor and the double descent effect, demonstrating that the perceived conflict between the two is, in fact, illusory. While the marginal likelihood in Bayesian model selection tends to favor simpler models, the conditional prior can still yield favorable risk properties even in highly over-parameterized regimes.
Third, we have established a comprehensive theoretical foundation for these ideas, drawing on Dawid’s model comparison theory, the Dickey-Savage results, and connections to generalized ridge regression and shrinkage methods. These theoretical developments provide a rigorous basis for understanding the interplay between prior specification, model complexity, and predictive risk.
Finally, we have illustrated how the Bayesian framework naturally extends to modern machine learning methods, such as neural networks, offering new insights into their generalization properties. This connection underscores the broad applicability of the Bayesian approach to contemporary challenges in statistical learning.
There are several important implications for cross-validation and the choice of regularization. This highlights the trade-off between training and test error in neural networks and other over-parameterized models. The Bayesian approach suggests that instead of relying solely on cross-validation for model selection, practitioners should also consider the implicit regularization provided by appropriate prior specifications.
There are many areas for future research. In higher dimensions, see the discussion in Polson et al. [2017] where one can find an ”island of good estimators.” Counterintuitively, this can be in the opposite direction of the classical bias-variance trade-off. Variance can be minimized by appealing to exchangeability results due to Kingman Kingman [1978] for exchangeable variables. See Amit and Geman [1997] for the discrete classification case and minimization of entropy of misclassification.
Fully eliciting priors and understanding their relationship with risk characteristics are still active areas of research. The connection between global-local shrinkage priors and double descent deserves further investigation, particularly in the context of deep learning where the number of parameters can be orders of magnitude larger than the sample size.
Another promising direction is the development of adaptive prior specifications that can automatically adjust to the data structure, potentially leading to more robust double descent behavior across different problem domains.